OpenStudy (anonymous):

A recipe calls for 2 1/3 cups of flour and 1 1/4 cups of sugar. If the recipe is tripled, how much flour and sugar will be needed?

OpenStudy (whpalmer4):

3 times as much :-)

OpenStudy (mathstudent55):

Convert each mixed numeral into an improper fraction, then multiply by 3/1.

OpenStudy (whpalmer4):

You could approach this at least 2 different ways: 1) addition 2) multiplication Addition: to make 3 batches of the recipe, add each amount 3 times. For example the flour would be 2 1/3 + 2 1/3 + 2 1/3 = 6 3/3 or 7 cups. Multiplication: multiply each amount by 3. 3*2 1/3 = 3*2 + 3*1/3 = 6+1 = 7

OpenStudy (whpalmer4):

What do you get for the amount of sugar?

OpenStudy (whpalmer4):

not that the problem asks, but it's generally preferable when baking to measure flour by weight...